OpenStudy (anonymous):

The square root of 5 more than twice a number is 7. Find the number

OpenStudy (anonymous):

22

OpenStudy (anonymous):

\[\sqrt {2x+5} = 7\] squaring both sides \[2x+ 5 = 7^2 \] \[2x+ 5 = 49\] \[2x= 49-5\] \[2x= 44\] \[x= \frac{44}{2}\] \[x = 22\]